# Resilience for Regular Path Queries: Towards a Complexity Classification

### v1

@@ 5.3. Implications of Hardness for Four-Legged Languages
We say that a language $L$ is star-free if there exists a $k>0$ such that, for every $\rho,\sigma,\tau\in\Sigma^{*}$, for every $m\geq k$, we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{m}\tau\in L$. Star-free languages have many equivalent characterizations (see, e.g., (mpri)). We can show that being star-free is a necessary condition for a reduced regular language to enjoy tractable resilience, because non-star-free languages are four-legged and hence resilience for them is NP-hard by Section 5.2:
@@ 6. Hardness of Finite Languages Having Repeated Letters
We now explain how Section 6.1 relates to the languages discussed so far. We first show a sanity check: such languages are never local.
@@ 7. Additional Complexity Results
We first illustrate two cases of finite languages that are tractable but not local. One are those we call bipartite chain languages, such as $ab|bc$. For them, we can compute resilience via a variant of the product construction of Theorem 3.3, simply by “reversing” some matches. Another is a class of languages (including for instance $abc|be$) which we show tractable via submodular minimization. These are the only tractable case where do not know how to reduce to a MinCut problem. We conclude by showing examples of queries not covered by our previous hardness results but for which resilience can be shown intractable, and finally exhibit queries that remain unclassified.
@@
We first deal in a special way with words of length 00 and $1$ that are accepted by $A$. We first test in linear time like in the proof of Section 7.2 whether $A$ accepts $\epsilon$, and if so, add $\epsilon$ to the list of accepted words.
@@
Thus, this construction maintains the one-to-one correspondence between the contingency sets of $Q_{L}$ for $D$ and the cuts of the network that have finite cost: selecting facts to remove in the contingency set corresponds to cutting edges in the network, and ensuring that the contingency set does not satisfy the query corresponds to ensuring that no source to target paths remain. Furthermore, this the sum of multiplicities of the facts in a contingency set is equal to the cost of the corresponding cut, so that minimum cuts correspond to minimum contingency sets. We can then, like in the case of local languages, simply solve the MinCut problem for this network in $\tilde{O}(|A|\times|D|^{2}\times|\Sigma|^{2})$, concluding the proof. ∎
@@ 8. Conclusion and Future Work
We have studied the complexity of resilience for RPQs, towards the goal of characterizing the languages for which the problem is tractable. We study bag and set semantics, but our results show no difference between the two settings, unlike for CQs (makhija2024unified). We leave open the question of whether there are languages $L$, regular or not, for which $\mathtt{RES}_{\mathrm{set}}(L)$ is PTIME but $\mathtt{RES}_{\mathrm{bag}}(L)$ is NP-hard.
Our work does not fully characterize the tractability boundary and leaves some cases open, like $abcd|be$ or $abc|bcd$ or $abc|bef$ or $ax^{*}b|xd$. At a high level, we still do not understand the non-local languages that are not four-legged, i.e., non-letter-Cartesian languages featuring counterexamples where some of the “legs” are empty. We also do not know if the submodularity approach is needed for queries like $abc|be$, or whether tractability can be obtained in easier ways. We also do not know how far the submodularity argument from Section 7.4 can extend to cover other queries.
Another open question concerns non-Boolean RPQs, i.e., where endpoints are fixed (or given as input). For instance, the RPQ $aa$ then seems tractable as possible contingency sets have a simple structure; but longer RPQs may stay hard. This problem may relate to length-bounded cuts (itai1982complexity; 10.1145/1868237.1868241), or multicuts (garg1994multiway) and skew multicuts (kratsch2015fixed). We also do not understand how our results relate to VCSPs (bodirsky2024complexity), e.g., whether classes of VCSPs with better algorithms (in combined complexity, or in terms of the polynomial degree) can imply tractable algorithms for some RPQ classes; or whether VCSP dichotomies hold in set semantics (or in bag semantics with weights restricted to $0,1,\infty$).
@@ 9. Acknowledgments
The authors are grateful to Charles Paperman for help about local languages. Part of this work was conducted while the authors were visiting the Simons Institute for the Theory of Computing. Amarilli was partially supported by the ANR project EQUUS ANR-19-CE48-0019, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 431183758, and by the ANR project ANR-18-CE23-0003-02 (“CQFD”). Gatterbauer and Makhija were partially supported by the National Science Foundation (NSF) under award number IIS-1762268.

### v2

@@ 5.3. Implications of Hardness for Four-Legged Languages
5.4. Reduced Languages of Star-Free Languages are Star-Free
We formally show the following result stated in the main text:
{claim}
Let $L$ be a star-free regular language. Then $\mathtt{reduce}(L)$ is star-free.
Of course, the converse statement is not true, e.g., $(aa)^{*}$ is not star-free but its reduced language $\epsilon$ is star-free. Let us prove the result:
Proof of Section 5.4.
We use an alternative characterization of star-free languages (mpri): they are those languages that can be constructed inductively by applying the operations of concatenation and the Boolean operations (including complementation) on base languages consisting of the empty language, the singleton language containing only the empty word, and the singleton languages formed of single-letter words for each letter of the alphabet $\Sigma$.
We then use the fact that, for any language $L$, we can express $\mathtt{reduce}(L)$ as follows:
$\mathtt{reduce}(L)=L\setminus(\Sigma^{+}L\Sigma^{*}\cup\Sigma^{*}L\Sigma^{+})$
Note that $\Sigma^{*}$ is the complement of the empty set, and $\Sigma^{+}$ is the concatenation of $\Sigma^{*}$ with the union of all letters of $\Sigma$, so both are star-free languages. Hence, if $L$ is a star-free language, then thanks to the fact that star-free languages are closed under concatenation and Boolean operations, so is $\mathtt{reduce}(L)$, which concludes the proof. ∎
@@ 5.4. Reduced Languages of Star-Free Languages are Star-Free
We say that a regular language $L$ is star-free if there exists a $k>0$ such that, for every $\rho,\sigma,\tau\in\Sigma^{*}$, for every $m\geq k$, we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{m}\tau\in L$. Star-free languages have many equivalent characterizations (see, e.g., (mpri)), and if a language $L$ is star-free then so is $\mathtt{reduce}(L)$ (see Section 5.4). We can show that being star-free is a necessary condition to enjoy tractable resilience, because non-star-free reduced regular languages are four-legged and hence resilience for them is NP-hard by Section 5.2:
@@ 6. Hardness of Finite Languages Having Repeated Letters
We now explain how Section 6.1 relates to the languages discussed so far. We first perform a sanity check: such languages are never local.
@@ 7. Additional Complexity Results
We first illustrate two cases of finite languages that are tractable but not local. One are those we call bipartite chain languages, such as $ab|bc$. For them, we can compute resilience via a variant of the product construction of Theorem 3.3, simply by “reversing” some matches. Another is a class of languages (including for instance $abc|be$) which we show tractable via submodular minimization. These are the only tractable case where we do not know how to reduce to a MinCut problem. We conclude by showing examples of queries not covered by our previous hardness results but for which resilience can be shown intractable, and finally exhibit queries that remain unclassified.
@@
We first deal in a special way with words of length 00 and $1$ that are accepted by $A$. We first test in linear time as in the proof of Section 7.2 whether $A$ accepts $\epsilon$, and if so, add $\epsilon$ to the list of accepted words.
@@
Thus, this construction maintains the one-to-one correspondence between the contingency sets of $Q_{L}$ for $D$ and the cuts of the network that have finite cost: selecting facts to remove in the contingency set corresponds to cutting edges in the network, and ensuring that the contingency set does not satisfy the query corresponds to ensuring that no source to target paths remain. Furthermore, this the sum of multiplicities of the facts in a contingency set is equal to the cost of the corresponding cut, so that minimum cuts correspond to minimum contingency sets. We can then, as in the case of local languages, simply solve the MinCut problem for this network in $\tilde{O}(|A|\times|D|^{2}\times|\Sigma|^{2})$, concluding the proof. ∎
@@ 8. Conclusion and Future Work
We have studied the complexity of resilience for RPQs, towards the goal of characterizing the languages for which the problem is tractable. We study bag and set semantics, but our results show no difference between the two settings, unlike for CQs (makhija2024unified). We leave open the question of whether there are languages $L$, regular or not, for which $\mathtt{RES}_{\mathrm{set}}(L)$ is PTIME but $\mathtt{RES}_{\mathrm{bag}}(L)$ is NP-hard. We also leave open the question of classifying the complexity of resilience in intermediate settings, e.g., bag semantics where only weights $1$, and $\infty$ are allowed; or the setting where some relations are indicated as exogenous in the signature, amounting to giving a weight of $\infty$ to all their facts.
Our work does not fully characterize the tractability boundary and leaves some cases open. At a high level, we still do not understand the non-local languages that are not four-legged, i.e., non-letter-Cartesian languages featuring counterexamples where some of the “legs” are empty. We also do not know if the submodularity approach is needed for queries like $abc|be$, or whether tractability can be obtained in easier ways. [^6] After we had submitted this paper for review, an argument was pointed out by Martín Muñoz (cstheory_martin) implying that Section 7.4 can be shown using a flow encoding instead of submodular function minimization. Further, his argument seems to imply that the tractability of resilience extends to more queries, such as $abcd|be$. We leave the exploration of these consequences to a future version of the paper.
Another open question concerns non-Boolean RPQs, i.e., where endpoints are fixed (or given as input). For instance, the RPQ $aa$ then seems tractable as possible contingency sets have a simple structure; but longer RPQs may stay hard. This problem may relate to length-bounded cuts (itai1982complexity; 10.1145/1868237.1868241), or multicuts (garg1994multiway) and skew multicuts (kratsch2015fixed). Note that this non-Boolean setting corresponds to the problem of deletion propagation with source side effects (buneman2002propagation), which was studied before resilience: this problem asks about finding a minimum contingency set to remove a specific query answer. The case of removing multiple tuples has also been investigated (kimelfeld2013multituple). We do not know whether a dichotomy in one of these settings implies a dichotomy for our problem, or vice-versa. We also do not understand how our results relate to VCSPs (bodirsky2024complexity), e.g., whether classes of VCSPs with better algorithms (in combined complexity, or in terms of the polynomial degree) can imply tractable algorithms for some RPQ classes; or whether VCSP dichotomies hold in set semantics (or in bag semantics with weights restricted to $1,\infty$).
@@ 9. Acknowledgments
The authors are grateful to Charles Paperman for help about local languages, and to the reviewers for their valuable feedback. Part of this work was conducted while the authors were visiting the Simons Institute for the Theory of Computing. Amarilli was partially supported by the ANR project EQUUS ANR-19-CE48-0019, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 431183758, and by the ANR project ANR-18-CE23-0003-02 (“CQFD”). Gatterbauer and Makhija were partially supported by the National Science Foundation (NSF) under award number IIS-1762268.